The small piston of a hydrolic lifts has an area of 0.2m^2 a car weighing 1.2*10^4N sit on a rack mounted on the large piston. The large piston has an area of 0.9m^2 how large a force must be applied to the small piston to support the car
1
Expert's answer
2018-12-06T13:21:11-0500

Answer on Question 83637, Physics, Other

Question:

The small piston of hydraulic lift has an area of

"0.2 m^2"

. A car weighing

"1.2*10^4 N"

sits on a rack mounted on the large piston. The large piston has an area of

"0.9 m^2"

. How large a force must be applied to the small piston to support the car?

Solution:

By the hydraulic press formula, we have:

"F_1\/A_1=F_2\/A_2"

,

here,

"A_1"

is the area of the small piston,

"A_2"

is the area of the large piston,

"F_1"

is the force applied to the small piston,

"F_2"

is the force applied to the large piston.

From this formula we can find the force applied to the small piston:

"F_1=F_2*(A_1\/A_2)=1.2*10^4 N*(0.2 m^2\/0.9 m^2)=2.7*10^3 N"

.

Answer:

"F_1=2.7*10^3 N"

.